OpenStudy (bigdog):

solve -3^2+2^0+27^(-2/3) so -3^2 is 9 and 2^0 is 1 which 9+1=10 then 27^-2/3= 1/27^(2/3) where do i go from there?

OpenStudy (saifoo.khan):

when ever you have a negative exponent, you have to flip the fraction.

OpenStudy (bigdog):

i did that with the 27 its now 1 over 27 does the 2/3 make the 27 1/9?

myininaya (myininaya):

(-3)^2=9 -3^2=-9 2^0=1 -3^2+2^0+27^(-2/3) -9+1+1/27^(2/3) -8+1/(3*3) -8+1/9 -8*9/9+1/9 -71/9

OpenStudy (bigdog):

thanks that helped alot becaues i was getting lost

OpenStudy (anonymous):

Ok your right so far with 1/27 ^(2/3) when you flipped the negative exponent, now you do 3√27 = 3 and take the 3, and do 3^2 which is 9, then you should be able to figure from there..

myininaya (myininaya):

-3^2 is different from (-3)^2